OpenStudy (anonymous):

I need help on Graphing Ellipses x^2/9 + y^2/1 = 1

OpenStudy (anonymous):

if the equation is in the form: \[\frac {(x-h)^2}{a^2}+\frac{(y-k)^2}{b^2}=1\] then it is easily graphed. the center is (h, k) a is the distance from your center to the endpoints of your major axis (vertices) b is the distance from your center to the endpoints of you minor axis (co-vertices) c is the distance from your center to the foci.

OpenStudy (anonymous):

your equation can be written in that form as: (x-0)^2/3^2 + (y-0)^2/1^2 = 1 so your center is at the origin; a = 3 b = 1 and here's the graph: